Consider the balanced chemical equation shown below.

4NH3(g) + 5O2(g) —> 4NO(g) + 6H2O(g)

In a certain experiment, 4.699 g of NH3(g) reacts with 1.850 g of O2(g).

A) Which is the limiting reactant?

B) How many grams of NO(g) form?

C) How many grams of H2O(g) form?

D) How many grams of the excess reactant remains after the limiting reactant is completely consumed?


1
Expert's answer
2021-02-22T06:09:17-0500

"4NH_{3}(g) + 5O_{2}(g)\rightarrow4NO(g)+6
H_{2}O(g)"


(a) According to stoichiometry,

5 moles of O2 reacts with 4 moles of NH3

160 g of O2 reacts with 68 g of NH3

1.850 g of O2 reacts with = "\dfrac{68}{160}\times1.850" g of NH3

= "0.78625" g of NH3

Since mass of O2 is fully consumed in the reaction and 3.91275 g of NH3 is in excess

"\therefore" O2 is the limiting reactant.


(b) 5 moles of O2 reacts to form 4 moles of NO

160 g of O2 react to form 132 g of NO

"\therefore" 1.850 g of O2 react to form = "\dfrac{132}{160}\times 1.850" g of NO

= "1.52625" g of NO


(c) Similarly mass of H2O can be found using stoichiometry,

Mass of H2O formed = "\dfrac{108}{160}\times1.850"

= "1.24875" g

(d) Mass of NH3(excess reactant) left = Total mass - consumed mass

= "4.699-0.78625"

= "3.91275" g